# Subgradient Convergence Implies Subdifferential Convergence on Weakly Convex Functions: With Uniform Rates Guarantees

### v1

@@ 2.4 Measurability Issues
This subsection can be mostly skipped on a first read, but it is essential for addressing measurability issues that arise when considering the supremum of random functions over an uncountable index set [Bil86]. This is important because we are interested in the supremum of the Hausdorff distance between stochastic subdifferential mappings over a given domain.
To address these measurability issues, we use the concept of outer measure $\mathbb{P}^{*}$ and inner measure $\mathbb{P}_{*}$ frequently adopted in empirical process theory [vdVW96, Section 1.2-5]. Given a probability space $(\Xi,\mathscr{G},\mathbb{P})$, every subset $B\subseteq\Xi$ is assigned with an outer measure:
@@ 3.1 Uniform Subdifferential Bounds via Selections
In this section, we introduce Theorem 3.1, a general technique for obtaining uniform bounds on the subdifferentials between two real-valued, locally weakly convex functions. The theorem states that the supremum of the Hausdorff distance between their subdifferentials over any open set is upper bounded by the supremum of the norm differences between any selected subgradient mappings.
Theorem 3.1 might initially seem surprising because at any single point $x$, the Hausdorff distance between the subdifferentials at $x$ can exceed the norm difference between the selected subgradients at $x$. To further appreciate this result, we will provide counterexamples in the remark to highlight why the assumption of $\mathcal{O}$ being an open set is crucial.
To the best of our knowledge, we are unaware of such a result documented in the literature. We note that in the (weakly) convex settings, graphical distance between subdifferentials under the Hausdorff metric is typically established through controlling the epi-distance of functions due to Attouch’s celebrated theorem [Att77, AB93, DD22]. See, e.g., [RW98, Theorem 12.35]. We shall make comparisons to clarify two distance metrics of subdifferential mappings in the second remark: the supremum of the Hausdorff distance between subdifferential mappings over $x$ in an open set $\mathcal{O}$, and the graphical distance between subdifferential mappings over the same open set $\mathcal{O}$.
Theorem 3.1.
@@
Our proof is structured into two parts. In the first part, we demonstrate that Theorem 3.1 holds under the assumption that $\mathcal{O}$ is open and convex, and that both $f_{1}$ and $f_{2}$ are convex on $\mathcal{O}$.
In the second part, we address the general case where $f_{1}$ and $f_{2}$ are locally weakly convex on an open set $\mathcal{O}$. We show that this scenario can be reduced to the convex case, thus concluding our proof of Theorem 3.1.
Part I.
In this part, we address the case where $f_{1}$ and $f_{2}$ are both real-valued convex functions on an open and convex set $\mathcal{O}$. Recall the definition of Hausdorff distance:
@@
where the interchangeability between inf and sup in the second identity is due to Sion’s minimax theorem [Sio58]. By substituting this identity back into equation (3.2), our goal reduces to prove:
@@
To prove this inequality, we invoke a perturbation argument standard in convex analysis. We start by taking a sequence $t_{n}\to 0$ as $n\to\infty$ with $t_{n}>0$ and $x+t_{n}v\in\mathcal{O}$ per $n\in\mathbb{N}$. Such a sequence $t_{n}$ exists because $\mathcal{O}$ is open. Next, we pick any cluster point $\tilde{y}_{2}$ of the sequence $\{g_{2}(x+t_{n}v)\}_{n\in\mathbb{N}}$. Given that $g_{2}(x+t_{n}v)\in\partial f_{2}(x+t_{n}v)$, and the subdifferential map $z\mapsto\partial f_{2}(z)$ is locally bounded [HUL96, Proposition 6.2.2], a cluster point $\tilde{y}_{2}$ necessarily exists in $\mathbb{R}^{d}$. Furthermore, because the subdifferential map $z\mapsto\partial f_{2}(z)$ is outer semicontinuous [HUL96, Theorem 6.2.4], $\tilde{y}_{2}$ must belong to the set $\partial f_{2}(x)$. Finally, by selecting a subsequence of $t_{n}$ if necessary, we can assume W.L.O.G. that the following limit holds:
@@
Given convexity of $f_{1}$ on the convex set $\mathcal{O}$, $y_{1}\in\partial f_{1}(x)$ and $g_{1}(x+t_{n}v)\in\partial f_{1}(x+t_{n}v)$, we obtain for every $n\in\mathbb{N}$:
@@
This shows that the desired inequality (3.4) holds for any $x\in\mathcal{O}$ and $y_{1}\in\partial f_{1}(x)$, and any $v$ with $\left\|{v}\right\|\leq 1$. Thus, we successfully establish the assertion of Theorem 3.1 in the convex settings.
Part II.
In the second part, we study the general case where $f_{1}$ and $f_{2}$ are both locally weakly convex on an open set $\mathcal{O}$, extending the established results for convex cases in Part I.
@@
Taking supremum over $z\in\mathcal{O}$, this observation directly leads to the conclusion of Theorem 3.1:
@@
Remark 3.2 (Necessity of $\mathcal{O}$ being an open set).
Theorem 3.1 would not hold if $\mathcal{O}$ were not an open set. Consider the convex functions $f_{1}(x)=|x|$ and $f_{2}(x)=2|x|$. Notably, the subdifferential sets are
@@
This illustrates the necessity for $\mathcal{O}$ to be an open set for Theorem 3.1 to hold. missing
Remark 3.3 (Comparison of Metrics on Subdifferential Mappings).
@@
where the Cartesian product $\mathcal{O}\times\mathbb{R}^{d}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}\cong\mathbb{R}^{2d}$ is equipped with the Euclidean distance. By definition, ${\rm d}_{1},{\rm d}_{2}$ are metrics on $\mathbb{S}_{\mathcal{O}}=\{g:\mathcal{O}\rightrightarrows\mathbb{R}^{d}:g=\partial f~{}\text{for some locally weakly convex $f:\mathcal{O}\to\mathbb{R}$}\}$. The metric ${\rm d}_{2}$ is known as the graphical distance, aligning well with applications of Attouch’s epigraphical convergence theorem [Att77], see also [DD22, Theorem 5.1].
@@ 3.2 Implications to Stochastic Weakly Convex Minimizations
Theorem 3.1 provides key insights into the uniform convergence of subdifferentials in stochastic minimizations. Theorem 3.4, which builds on top of Theorem 3.1, shows that for stochastic minimizing objectives with certain weak convexity requirements, the uniform convergence of subdifferential sets can be effectively understood by analyzing the uniform convergence of any pair of selections from the subdifferentials of both population and empirical objectives.
To formalize our results, we begin by introducing our setup. Let $f(\cdot,\cdot):\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ denote a real-valued random function. We are sampling $\xi\sim\mathbb{P}$ where $\mathbb{P}$ is supported on $\Xi$, assumed to be a Euclidean space. We assume $\mathbb{P}$ is a complete measure to avoid some measurability issues.
@@
In the above, $\xi_{1},\xi_{2},\ldots,\xi_{m}$ are i.i.d. samples drawn from the distribution $\mathbb{P}$, where $\mathbb{P}_{m}$ denotes the empirical distribution of $\xi_{1},\xi_{2},\ldots,\xi_{m}$. The function $R:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex function. The constraint set $X$ is a nonempty, closed, and convex set. We are interested in establishing an upper bound for the rate of convergence of subdifferentials of the empirical objective $\phi_{S}$ to the population objective $\phi$ over some subset of $X$, frequently modeled as $X\cap\mathcal{O}$ where $\mathcal{O}$ is an open set in $\mathbb{R}^{d}$. This openness requirement for $\mathcal{O}$ originates from Theorem 3.1 and, despite this condition, our result is relevant to a broad range of interesting applications.
@@
is closed and convex on the ball ${\rm cl}(\mathbb{B}(x;\epsilon(x)))$ and $\mathbb{E}[\lambda(x,\xi)]<\infty$.
Assumption B immediately implies the function $f_{S}(x)$ is locally weakly convex on $\mathcal{O}$ (with probability one). To see this, for each $x\in\mathcal{O}$,
@@
is convex on the open neighborhood $\mathbb{B}(x;\epsilon(x))$. Similarly, the function $f(x)=\mathbb{E}[f_{S}(x)]$ is locally weakly convex within $\mathcal{O}$.
Building on Theorem 3.1, Theorem 3.4 naturally follows.
Theorem 3.4.
@@
$\sup_{x\in\mathcal{O}\cap X}\mathbb{H}(\phi(x),\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|.$
@@
Given that $X$ is closed and convex, $\iota_{X}$ is subdifferentially regular at any $x\in X$ [RW98, Exercise 8.14]. The functions $f(x)$ and $f_{S}(x)$ are real-valued, closed and weakly convex on $\mathcal{O}$ by Assumptions A and B. Thus they are subdifferentially regular on $\mathcal{O}$ [RW98, Corollary 8.11], with $\partial^{\infty}f(x)=\partial^{\infty}f_{S}(x)=\{0\}$ for all $x\in\mathcal{O}$. Similarly, the real-valued, closed, and convex function $r$, is subdifferentially regular at $x\in\mathcal{O}$ with $\partial^{\infty}R(x)=\{0\}$.
These regularity properties of the functions enable the application of basic subdifferential calculus [RW98, Corollary 10.9]. Accordingly, the subdifferentials at $x\in X$ are given by:
@@
The last inequality is due to Theorem 3.1, acknowledging that, with probability one, $f$ and $f_{S}$ are real-valued and weakly-convex on the open set $\mathcal{O}$. ∎
@@ 3.3 A Class of Stochastic Weakly Convex Minimizations
In the literature, it is known that under mild conditions, the convex composition $f(\cdot,\xi)=h(c(\cdot;\xi))$ satisfies weak-convexity requirements in Assumption B, resulting in locally weakly convex objectives $f(x)$ and $f_{S}(x)$ that fit into our diagrams in Section 3.2, see, e.g., [DR18, Claim 1] and [DD19]. We will revisit these conditions subsequently before we present our main results in Theorem 3.9.
We give examples of stochastic convex-composite objectives that appear in statistics, machine learning, imaging, and risk management. More examples can be found in [DD19, Section 2.1].
Example 3.5 (Robust Phase Retrieval).
Phase retrieval is a computational problem with applications across various fields, including imaging, X-ray crystallography, and speech processing. The (real-valued) phase retrieval seeks to detect a point $x$ satisfying $|\langle a_{i},x\rangle|^{2}=b_{i}$ where $a_{i}\in\mathbb{R}^{d}$, and $b_{i}\in\mathbb{R}$ for $i=1,2,\ldots,m$. We can choose $\xi=(a,b)\in\mathbb{R}^{d}\times\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=(a^{T}x)-b$, and $X=\mathbb{R}^{d}$, in which case the form (3.9) gives an exact penalty formulation for solving the collection of quadratic equations, which yields strong statistical recovery and robustness guarantees [EM14, DR19, DDP20], among other nonconvex formulations [CLS15, WGE17, SQW18].
Example 3.6 (Robust Matrix Sensing).
This problem can be viewed as a variant of phase retrieval. Let $A_{1},A_{2},\ldots,A_{m}\in\mathbb{R}^{D\times D}$ be measurement matrices. Given the measurement $b_{i}=\langle A_{i},M_{\sharp}\rangle+\eta_{i}\in\mathbb{R}$, where $M_{\sharp}\in\mathbb{R}^{D\times D}$ is the true matrix, and $\eta_{i}\in\mathbb{R}$ is the noise corruption, the goal is to recover a low-rank approximation of $M_{\sharp}$, which is modeled through $XX^{T}$ where $X\in\mathbb{R}^{D\times r_{0}}$ where $1\leq r_{0}\leq D$. Recently, there has been a series of efforts showing that the following potential function has strong stability guarantees under certain assumptions on $A$ and $\eta$ [LZMCSV20, CCD+21, DJC+21]:
@@
Example 3.7 (Conditional Value-at-Risk).
Let $\ell(w,\xi)$ represent a decision rule parameterized by $w\in\mathbb{R}^{d}$ for a data point $\xi$, where $\xi\sim\mathbb{P}$. Instead of minimizing the expected value $\mathbb{E}_{\xi\sim\mathbb{P}}[\ell(w,\xi)]$, it is often preferable to minimize the conditional expectation of the random variable $\ell(w,\cdot)$ over its $\alpha$-tail, where $\alpha\in(0,1)$ is given. This quantity is termed the Conditional Value-at-Risk (cVaR) [SDR21, Section 6]. Remarkably, a seminal work shows that minimizing cVaR can be formulated as [RU00]:
@@
We now revisit the conditions on the convex function $h$ and smooth function $c(x,\xi)$ that guarantees the weak-convexity assumptions for the composition $f(x,\xi)=h(c(x,\xi))$, ensuring the regularity required for subdifferential calculus rules. We take the following conditions—local Lipschitzian and integrability conditions on $h,c$—from [DR18, Section 2.2].
@@
Given Assumptions A’ and B’, the following basic property on the stochastic convex-composite objectives follows from [DR18, Claim 1+Lemma 3.6]. We use $\nabla c(x;\xi)$ to denote the gradient of a smooth function $c$ with respect to $x$ for a given $\xi$.
Lemma 3.8.
@@
For ease of reference in future discussions, we establish a result that stems directly from Theorem 3.4 and Lemma 3.8 for stochastic convex-composite minimizations.
Theorem 3.9.
@@
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\phi(x),\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|.$

### v2

@@ 2.4 Measurability Issues
This subsection can be mostly skipped on a first read, but it is essential for addressing measurability issues that arise when considering the supremum of random functions over an uncountable index set [30]. This is important because we are interested in the supremum of the Hausdorff distance between stochastic subdifferential mappings over a given domain.
To address these measurability issues, we use the concept of outer measure $\mathbb{P}^{*}$ and inner measure $\mathbb{P}_{*}$ frequently adopted in empirical process theory [31, Section 1.2-5]. Given a probability space $(\Xi,\mathscr{G},\mathbb{P})$, every subset $B\subseteq\Xi$ is assigned with an outer measure:
@@ 3.1 Uniform Subdifferential Bounds via Selections
In this section, we introduce Theorem 1, a general technique for obtaining uniform bounds on the subdifferentials between two real-valued, locally weakly convex functions. The theorem states that the supremum of the Hausdorff distance between their subdifferentials over any open set is upper bounded by the supremum of the norm differences between any selected subgradient mappings.
Theorem 1 might initially seem surprising because at any single point $x$, the Hausdorff distance between the subdifferentials at $x$ can exceed the norm difference between the selected subgradients. To further appreciate this result, we will provide counterexamples in the remark to highlight why the assumption of $\mathcal{O}$ being an open set is crucial.
We note that in the (weakly) convex settings, graphical distance between subdifferentials under the Hausdorff metric is typically established through controlling the epi-distance of functions due to Attouch’s celebrated theorem [15, 32, 16]. See, e.g., [17, Theorem 12.35]. We shall make comparisons to clarify two distance metrics of subdifferential mappings in the second remark: the supremum of the Hausdorff distance between subdifferential mappings over $x$ in an open set $\mathcal{O}$, and the graphical distance between subdifferential mappings over the same open set $\mathcal{O}$.
Theorem 1.
@@
Our proof is structured into two parts. In the first part, we demonstrate that Theorem 1 holds under the assumption that $\mathcal{O}$ is open and convex, and that both $f_{1}$ and $f_{2}$ are convex on $\mathcal{O}$.
In the second part, we address the general case where $f_{1}$ and $f_{2}$ are locally weakly convex on an open set $\mathcal{O}$. We show that this scenario can be reduced to the convex case, thus concluding our proof of Theorem 1.
Part I. In this part, we address the case where $f_{1}$ and $f_{2}$ are both real-valued convex functions on an open and convex set $\mathcal{O}$. Recall the definition of Hausdorff distance:
@@
where the interchangeability between inf and sup in the second identity is due to Sion’s minimax theorem [33]. By substituting this identity back into equation (3.2), our goal reduces to prove:
@@
To prove this inequality, we invoke a perturbation argument standard in convex analysis. We start by taking a sequence $t_{n}\to 0$ as $n\to\infty$ with $t_{n}>0$ and $x+t_{n}v\in\mathcal{O}$ per $n\in\mathbb{N}$. Such a sequence $t_{n}$ exists because $\mathcal{O}$ is open. Next, we pick any cluster point $\tilde{y}_{2}$ of the sequence $\{g_{2}(x+t_{n}v)\}_{n\in\mathbb{N}}$. Given that $g_{2}(x+t_{n}v)\in\partial f_{2}(x+t_{n}v)$, and the subdifferential map $z\mapsto\partial f_{2}(z)$ is locally bounded [34, Proposition 6.2.2], a cluster point $\tilde{y}_{2}$ necessarily exists in $\mathbb{R}^{d}$. Furthermore, because the subdifferential map $z\mapsto\partial f_{2}(z)$ is outer semicontinuous [34, Theorem 6.2.4], $\tilde{y}_{2}$ must belong to the set $\partial f_{2}(x)$. Finally, by selecting a subsequence of $t_{n}$ if necessary, we can assume W.L.O.G. that the following limit holds:
@@
Given convexity of $f_{1}$, $y_{1}\in\partial f_{1}(x)$ and $g_{1}(x+t_{n}v)\in\partial f_{1}(x+t_{n}v)$, we obtain for every $n\in\mathbb{N}$:
@@
This shows that the desired inequality (3.4) holds for any $x\in\mathcal{O}$ and $y_{1}\in\partial f_{1}(x)$, and any $v$ with $\left\|{v}\right\|\leq 1$. Thus, we successfully establish the assertion of Theorem 1 in the convex settings.
Part II. In the second part, we study the general case where $f_{1}$ and $f_{2}$ are both locally weakly convex on an open set $\mathcal{O}$, extending the established results for convex functions in Part I.
@@
Taking supremum over $z\in\mathcal{O}$, this observation directly leads to the conclusion of Theorem 1:
@@
Remark A (Necessity of $\mathcal{O}$ being an open set).
Theorem 1 would not hold if $\mathcal{O}$ were not an open set. Consider the convex functions $f_{1}(x)=|x|$ and $f_{2}(x)=2|x|$. Notably, the subdifferential sets are
@@
This illustrates the necessity for $\mathcal{O}$ to be an open set for Theorem 1 to hold. missing
Remark B (Comparison of Metrics on Subdifferential Mappings).
@@
where the Cartesian product $\mathcal{O}\times\mathbb{R}^{d}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}\cong\mathbb{R}^{2d}$ is equipped with the Euclidean distance. By definition, ${\rm d}_{1},{\rm d}_{2}$ are metrics on $\mathbb{S}_{\mathcal{O}}=\{g:\mathcal{O}\rightrightarrows\mathbb{R}^{d}:g=\partial f~{}\text{for some locally weakly convex $f:\mathcal{O}\to\mathbb{R}$}\}$. The metric ${\rm d}_{2}$ is known as the graphical distance, aligning well with applications of Attouch’s epigraphical convergence theorem [15], see also [16, Theorem 5.1].
@@ 3.2 Implications to Stochastic Weakly Convex Minimizations
Theorem 1 provides key insights into the uniform convergence of subdifferentials in stochastic minimizations. Theorem 2, which builds on top of Theorem 1, shows that for stochastic minimizing objectives with certain weak convexity requirements, the uniform convergence of subdifferential sets can be effectively understood by analyzing the uniform convergence of any pair of selections from the subdifferentials of both population and empirical objectives.
To formalize our results, we begin by introducing our setup. Let $f(\cdot,\cdot):\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ denote a real-valued random function. We are sampling $\xi\sim\mathbb{P}$ where $\mathbb{P}$ is supported on $\Xi$, assumed to be a Euclidean space.
@@
In the above, $\xi_{1},\xi_{2},\ldots,\xi_{m}$ are i.i.d. samples drawn from the distribution $\mathbb{P}$, where $\mathbb{P}_{m}$ denotes the empirical distribution of $\xi_{1},\xi_{2},\ldots,\xi_{m}$. Formally, we start with a single random variable $\xi$ defined on a probability space $(\Xi,\mathcal{G},\mathbb{P})$. Utilizing this space, we can construct the countable product space $(\Xi,\mathcal{G},\mathbb{P})^{\mathbb{N}}$ where the sequence of i.i.d. random variables $(\xi_{1},\xi_{2},\ldots,\xi_{m},\ldots)$ can be properly defined on this product probability space [30, Section 36], with each $\xi_{i}$ acting as a coordinate projection. In this product space, we then extend the sigma-algebra and the probability measure to obtain a complete measure, following standard procedures in measure theory [30], to avoid some cumbersome measurability issues [2, Section 7]. The function $R:\mathbb{R}^{d}\to\mathbb{R}$ is a real-valued, closed and convex function. The constraint set $X$ is a nonempty, closed, and convex set.
We are interested in establishing an upper bound for the rate of convergence of subdifferentials of the empirical objective $\phi_{S}$ to the population objective $\phi$ over some subset of $X$, frequently modeled as $X\cap\mathcal{O}$ where $\mathcal{O}$ is an open set in $\mathbb{R}^{d}$. This openness requirement for $\mathcal{O}$ originates from Theorem 1 and, despite this condition, our result is relevant to a broad range of interesting applications.
@@
is convex on the open ball $\mathbb{B}(x;\epsilon(x))$ and $\mathbb{E}_{\xi\sim\mathbb{P}}[\lambda(x,\xi)]<\infty$.
Lemma 1.
Let Assumptions A and B hold. Then with probability one, the function $f_{S}(x)$ is locally weakly convex on the open set $\mathcal{O}$. Also, the function $f(x)=\mathbb{E}[f_{S}(x)]$ is locally weakly convex on $\mathcal{O}$.
Proof.
By Assumption B, for each $x\in\mathcal{O}$, there exists a radius $\epsilon(x)>0$ and $\lambda(x,\xi)\geq 0$ such that
$y\mapsto f(y,\xi)+\frac{\lambda(x,\xi)}{2}\left\|{y}\right\|^{2}$
is convex on the open ball $\mathbb{B}(x;\epsilon(x))$. The collection of these balls, $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}}$ forms an open cover of $\mathcal{O}$. By Lindelöf’s covering Theorem, see, e.g., [35, Page 50], and noting that $\mathbb{R}^{d}$ with its norm topology is second countable, there is a countable subcover $\{\mathbb{B}(x;\epsilon(x))\}_{x\in\mathcal{O}^{o}}$ of $\mathcal{O}$, where $\mathcal{O}^{o}$ is a countable subset of $\mathcal{O}$. In particular, $\mathcal{O}$ obeys the inclusion:
$\mathcal{O}\subseteq\bigcup_{x\in\mathcal{O}^{o}}\mathbb{B}(x;\epsilon(x)).$
For each $x\in\mathcal{O}^{o}$, the function
@@
is convex on the open neighborhood $\mathbb{B}(x;\epsilon(x))$. Given that $\mathbb{E}_{\xi\sim\mathbb{P}}[\lambda(x,\xi)]<\infty$ holds, for each $x\in\mathcal{O}^{o}$ it follows that $\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\lambda(x,\xi)]<\infty$ with probability one. Using a union bound over the countable set $\mathcal{O}^{o}$ this implies that with probability one, $\mathbb{E}_{\xi\sim\mathbb{P}_{m}}[\lambda(x,\xi)]<\infty$ simultaneously holds for every $x\in\mathcal{O}^{o}$. Thus, with probability one $f_{S}(x)$ is locally weakly convex on the set $\mathcal{O}$, as $\mathcal{O}$ is included in the countable union of these open balls: $\mathcal{O}\subseteq\bigcup_{x\in\mathcal{O}^{o}}\mathbb{B}(x;\epsilon(x))$. A similar reasoning yields that the function $f(x)=\mathbb{E}[f_{S}(x)]$ is locally weakly convex within the open set $\mathcal{O}$. ∎
Building on Theorem 1 and Lemma 1 then Theorem 2 naturally follows.
Theorem 2.
@@
$\sup_{x\in\mathcal{O}\cap X}\mathbb{H}(\partial\phi(x),\partial\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|.$
@@
Given that $X$ is closed and convex, $\iota_{X}$ is subdifferentially regular at any $x\in X$ [17, Exercise 8.14]. The functions $f(x)$ and $f_{S}(x)$ are real-valued, closed and weakly convex on $\mathcal{O}$ by Assumptions A and B. Thus they are subdifferentially regular on $\mathcal{O}$ [17, Corollary 8.11], with $\partial^{\infty}f(x)=\partial^{\infty}f_{S}(x)=\{0\}$ for all $x\in\mathcal{O}$. Similarly, the real-valued, closed, and convex function $r$, is subdifferentially regular at $x\in\mathcal{O}$ with $\partial^{\infty}R(x)=\{0\}$.
These regularity properties of the functions enable the application of basic subdifferential calculus [17, Corollary 10.9]. Accordingly, the subdifferentials at $x\in X$ are given by:
@@
The last inequality is due to Theorem 1, acknowledging that, with probability one, $f$ and $f_{S}$ are real-valued and weakly-convex on the open set $\mathcal{O}$ by Lemma 1. This inequality, which involves the supremum over uncountable random variables, represents a measurable event with probability one due to the completeness of the probability measure on the product space (see the setup on the probability space at the beginning of this subsection). ∎
@@ 3.3 A Class of Stochastic Weakly Convex Minimizations
In the literature, it is known that under mild conditions, the convex composition $f(\cdot,\xi)=h(c(\cdot;\xi))$ satisfies weak-convexity requirements in Assumption B, resulting in locally weakly convex objectives $f(x)$ and $f_{S}(x)$ that fit into our diagrams in Section 3.2, see, e.g., [7, Claim 1] and [36]. We will revisit these conditions subsequently before we present our main results in Theorem 3.
We give examples of stochastic convex-composite objectives that appear in statistics, machine learning, imaging, and risk management. More examples can be found in [36, Section 2.1].
Example 1 (Robust Phase Retrieval).
Phase retrieval is a computational problem with applications across various fields, including imaging, X-ray crystallography, and speech processing. The (real-valued) phase retrieval seeks to detect a point $x$ satisfying $|\langle a_{i},x\rangle|^{2}=b_{i}$ where $a_{i}\in\mathbb{R}^{d}$, and $b_{i}\in\mathbb{R}$ for $i=1,2,\ldots,m$. We can choose $\xi=(a,b)\in\mathbb{R}^{d}\times\mathbb{R}$, $h(z)=|z|$ and $c(x;\xi)=(a^{T}x)-b$, and $X=\mathbb{R}^{d}$, in which case the form (3.9) gives an exact penalty formulation for solving the collection of quadratic equations, which yields strong statistical recovery and robustness guarantees [37, 8, 14], among other nonconvex formulations [38, 39, 40].
Example 2 (Robust Matrix Sensing).
This problem can be viewed as a variant of phase retrieval. Let $A_{1},A_{2},\ldots,A_{m}\in\mathbb{R}^{D\times D}$ be measurement matrices. Given the measurement $b_{i}=\langle A_{i},M_{\sharp}\rangle+\eta_{i}\in\mathbb{R}$, where $M_{\sharp}\in\mathbb{R}^{D\times D}$ is the true matrix, and $\eta_{i}\in\mathbb{R}$ is the noise corruption, the goal is to recover a low-rank approximation of $M_{\sharp}$, which is modeled through $XX^{T}$ where $X\in\mathbb{R}^{D\times r_{0}}$ where $1\leq r_{0}\leq D$. Recently, there has been a series of efforts showing that the following potential function has strong stability guarantees under certain assumptions on $A$ and $\eta$ [10, 11, 41]:
@@
Example 3 (Conditional Value-at-Risk).
Let $\ell(w,\xi)$ represent a decision rule parameterized by $w\in\mathbb{R}^{d}$ for a data point $\xi$, where $\xi\sim\mathbb{P}$. Instead of minimizing the expected value $\mathbb{E}_{\xi\sim\mathbb{P}}[\ell(w,\xi)]$, it is often preferable to minimize the conditional expectation of the random variable $\ell(w,\cdot)$ over its $\alpha$-tail, where $\alpha\in(0,1)$ is given. This quantity is termed the Conditional Value-at-Risk (cVaR) [2, Section 6]. Remarkably, a seminal work shows that minimizing cVaR can be formulated as [12]:
@@
We now revisit the conditions on the convex function $h$ and smooth function $c(x,\xi)$ that guarantees the weak-convexity assumptions for the composition $f(x,\xi)=h(c(x,\xi))$, ensuring the regularity required for subdifferential calculus rules. We take the following conditions—local Lipschitzian and integrability conditions on $h,c$—from [7, Section 2.2].
@@
Given Assumptions A’ and B’, the following basic property on the stochastic convex-composite objectives follows from [7, Claim 1+Lemma 3.6]. We use $\nabla c(x;\xi)$ to denote the gradient of a smooth function $c$ with respect to $x$ for a given $\xi$.
Lemma 2.
@@
For ease of reference in future discussions, we establish a result that stems directly from Theorem 2 and Lemma 2 for stochastic convex-composite minimizations.
Theorem 3.
@@
$\sup_{x\in X\cap\mathcal{O}}\mathbb{H}(\partial\phi(x),\partial\phi_{S}(x))\leq\sup_{x\in\mathcal{O}}\left\|{G(x)-G_{S}(x)}\right\|.$